OpenStudy (anonymous):

@Arnab09 Factor 6x^2 + 5x – 4. A. (3x + 4)(2x – 1) B. (3x – 1)(3x + 4) C. (6x – 4)(x + 1) D. The polynomial is prime.

OpenStudy (anonymous):

6x^2+5x-4= 6x^2+8x-3x-4= 2x(3x+4)-1(3x+4)=(2x-1)(3x+4)